Question 6 [2]


State the mean Value Theorem.

1
Expert's answer
2022-03-10T18:13:41-0500

The theorem states that if "f" is a continuous function on the closed interval "[a,b]" and differentiable on the open interval "(a,b)," then there exists a point "c" in "(a,b)" such that the tangent at "c" is parallel to the secant line through the endpoints "{\displaystyle {\big (}a,f(a){\big )}}" and "{\displaystyle {\big (}b,f(b){\big )}}," that is,


"{\displaystyle f'(c)={\frac {f(b)-f(a)}{b-a}}.}"